A discriminatory movie monopolist sells the service to college students and lecturers. The demand function for students is 𝑄s = 200 βˆ’ 25𝑃 and the demand function for lecturers is


𝑄l = 400 βˆ’ 25P. Marginal cost is $5 per ticket.



(a) Which degree (type) of price discrimination the monopolist applied? Why?



(b) Calculate the amount of ticket the firm sells for students.



(c) Calculate the amount of ticket the firm sells for lecturers.




1
Expert's answer
2022-03-01T14:50:43-0500

a) Third-degree price discrimination- occurs when a company charges a different price to different consumer groups.

b) Amount of ticket for students

"\ud835\udc44_s = 200 \u2212 25\ud835\udc43"

P=8- 0.04Q

TR= Q(8- 0.04Q)= 8Q-0.04Q"^2"

MR= 8- 0.08Q

MC= 5

MR= MC

8- 0.08Q= 5

Q= 37.5

c) Amount of ticket for lecturers

MR= MC

"Q_s" =400 βˆ’ 25P

P= 16- 0.04Q

TR= 16Q- 0.04Q"^2"

MR= 16- 0.08Q

MR= MC

16- 0.08Q= 5

Q= 137.5